OpenStudy (czarluc):

sin3theta + sin2theta = 1?

OpenStudy (gorv):

u know formula of sin3 theta??

OpenStudy (czarluc):

wat is it? I only know the double angle ones

OpenStudy (gorv):

okayyy...so \[\sin3\theta=\sin(2\theta+\theta)\]

OpenStudy (gorv):

this u can expand ??

OpenStudy (czarluc):

uhm. general addition formula?

OpenStudy (gorv):

use sin(A+B) A=2 theta B=Theta

OpenStudy (gorv):

yeah can u expand it ??

OpenStudy (czarluc):

yes

OpenStudy (czarluc):

oh ok . then I would just do some algebraic manipulations.. thank you very much!

OpenStudy (gorv):

yeah ..u got it ..

OpenStudy (gorv):

welcome:)